OpenStudy (hashpipelike17):

4.012 in expanded. form

satellite73 (satellite73):

what does "expanded form" mean?

satellite73 (satellite73):

like maybe \[3.14=3\times 10^0+1\times 10^{-1}+4\times 10^{-2}\] or sommat?

OpenStudy (hashpipelike17):

When we write the number 521, what that number really means is that we have the total of 500 + 20 + 1. We've expanded the number to show the value of each of its digits. When we expand a number to show the value of each digit, we're writing that number in expanded form.

satellite73 (satellite73):

ok so it is the same only now \[4+\frac{0}{10}+\frac{1}{100}+\frac{2}{1000}\] i guess

OpenStudy (hashpipelike17):

ok thxs

satellite73 (satellite73):

yw